OpenStudy (anonymous):

sqrt(3+4)sqrt(3) simplified in radical form

sam (.sam.):

Add (3+4) then use \[\sqrt{a} \cdot \sqrt{b}=\sqrt{a \times b}\]

OpenStudy (anonymous):

sqrt 7sqrt 3?